Two masses are tied to the ends of a rope, and the
Answer #1

here

Fnet = Flarge - Fsmall = mlarge*g - msmall *g

Fnet = 20.5*9.81 - 14.7 * 9.81

Fnet = 56.89 N downward

then

a = Fnet / m1 + m2

a = 56.89 / 35.2 = 1.61 m/s^2